OpenStudy (anonymous):

If x+y=2, what is the minimum value of x^2+y^2?

OpenStudy (jiteshmeghwal9):

by squaring (x+y=2)\[\huge{(x+y)^2=(2)^2}\]\[\huge{x^2+y^2+2xy=4}\]\[\huge{x^2+y^2=4-2xy}\]

OpenStudy (anonymous):

So how do you find the minimum value?

OpenStudy (jiteshmeghwal9):

Is there given any value of xy ??/

OpenStudy (anonymous):

Nope.

OpenStudy (jiteshmeghwal9):

@ganeshie8 sir plz help :)

OpenStudy (raden):

this is a quadratic function problem... x+y=2 ---> y=2-x x^2+y^2 = x^2 + (2-x)^2 let k=x^2+y^2, so k=x^2 + (2-x)^2 k=x^2 + 4 - 4x + x^2 k=2x^2 - 4x + 4 k would be minimum when x=-b/2a=-(-4)/2*2 = 4/4 = 1 to find the minimum value of k, just plug x=1 k=2x^2 - 4x + 4 k=2(1)^2 - 4(1) + 4 k=2-4+4 k=2

OpenStudy (anonymous):

for minimum value the two value should be close as possible ,,so here it should be 1,1..so minimum value should be 2

OpenStudy (anonymous):

Where did the equation x=-b/2a come from?

OpenStudy (raden):

it is x-vertec of quadratic function y=ax^2+bx+c

OpenStudy (anonymous):

Ok, thank you.